OpenStudy (anonymous):

Solve: n4 - 2n2 - 8 = 0

OpenStudy (anonymous):

n4 - 2n2 - 8 = 0 n4 -4n2 +2n2 -8 = 0 n2(n2-4)+2(n2 -4) =0 (n2+2)(n2-4) = 0 now take it from here

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

welcome